# Generalized Rate Operator Quantum Jumps via Realization-Dependent Transformations

### v1

@@ V Control of the realizations
Figure 2: Eternally non-Markovian dynamics. Top: $z$ (blue, solid) and $x$ (green, dashed) components of the Bloch vectors. The thick solid lines are the exact results, the dots are obtained with the RO technique. In lighter shade, 5 realizations are shown. The unravelings are obtained using $N=10^{3}$ states. (a): $\phi_{1}=\phi_{1}^{\text{ub}}$ and only jumps to |$\ket{0}$; (b): $\phi_{1}=\phi_{1}^{\text{lb}}$ and only jumps to |$\ket{1}$; (c): jumps to $\ket{\pm}$. Bottom: characterization of the stochastic realizations for different choices of $\phi_{1}=\lambda\phi_{1}^{\text{lb}}+(1-\lambda)\lambda\phi_{1}^{\text{ub}}$, $0\leq\lambda\leq 1$. (d): Shannon entropy $H(\{p_{i}\})=-\sum_{i}p_{i}\log_{2}p_{i}$ for the probability distribution $\{p_{0},p_{1},p_{\psi}\}$ of the occupations of the states |$\ket{0},\ket{1},\ket{\psi_{\text{det}}(t)}$; (e): number of jumps, using $N=10^{4}$ states; (f): computational time (left axis, solid) and total number of jumps (right axis, dotted).
@@ V.1 Phase covariant dynamics
We now show that, as long as P-divisibility holds, it is always possible to have a positive unraveling using only three states: the initial state |$\ket{\psi_{\text{det}}(t)}$, deterministically evolving via (19), and the eigenstates |$\ket{0},\ket{1}$ of $\sigma_{z}$. The possibility of using such a small effective ensemble drastically simplifies the simulations. The RO is chosen such that |$\ket{\psi_{\text{det}}(t)}$ only jumps to |$\ket{0},\ket{1}$ and, after one jump has occurred, only jumps |$\ket{1}\leftrightarrow\ket{0}$ are present, without any deterministic evolution. Thus, the effective ensemble used for the simulations only contains three states. The possibility of considering such finite (and small) effective ensemble, significantly enhances the computational efficiency of this method, since one is not required to compute at each time-step the evolution of all states, but only needs to update the probability of occupation of such states. This fact is particularly interesting in comparison with the NMQJ method that, expect for some special cases, needs infinitely many states in the effective ensemble.
[…]
|$\ket{\Phi_{\psi}}=\alpha\left(2\gamma_{z}-\frac{\phi_{1}}{\sqrt{1-\lvert\alpha\rvert^{2}}}\right)\ket{0}+\phi_{1}\ket{1},$ (27)
where |$\ket{\psi}=\alpha\ket{0}+\sqrt{1-\lvert\alpha\rvert^{2}}\ket{1}$, with $\alpha$ that, without loss of generality, can be assumed to be real because of phase covariance, while $\phi_{1}$ which can be chosen freely inside a suitable time- and state-dependent interval $\phi_{1}\in[\phi_{1}^{\text{lb}},\phi_{1}^{\text{ub}}]$ (for the details, the definition of $\phi_{1}^{\text{lb}},\phi_{1}^{\text{ub}}$ and the proof of the positivity, see Appendix C). The freedom in choosing $\phi_{1}$ allows us to have different realizations for the unraveling: even if the post-jump states are fixed, it is possible to unravel the dynamics with different jump rates and deterministic evolutions. In particular, for $\phi_{1}=\phi_{1}^{\text{lb}}$, only jumps |$\ket{\psi}\mapsto\ket{1}$ are allowed, while for $\phi_{1}=\phi_{1}^{\text{ub}}$, only |$\ket{\psi}\mapsto\ket{0}$. On the other hand, for any $\phi_{1}=\lambda\phi_{1}^{\text{lb}}+(1-\lambda)\phi_{1}^{\text{ub}}$, $10<\lambda<$ 1, the unraveling remains positive, with jumps to both eigenstates of $\sigma_{z}$ and with the possibility of choosing different jump rates depending on $\lambda$.
@@ V.1.1 Eternally non-Markovian dynamics
The negativity of $\gamma_{z}$ at all times implies that such dynamics is CP-indivisible at all times, and therefore cannot be unraveled using the standard MCWF nor its non-Markovian generalization [30, 29] because of the negativity of the rate since the very beginning of time. However, it is possible to unravel it using the generalized RO. In addition, it is possible to realize qualitatively different realizations of the stochastic process. In Fig. 2 (a)-(b), we show unravelings obtained either with only jumps to |$\ket{0}$ or to |$\ket{1}$ by suitably choosing $\phi_{1}$. The code used for obtaining the simulations is available at [62]. Additionally, it can also be unraveled with jumps to |±=(|$\ket{\pm}=(\ket{0}\pm\ket{1})/\sqrt{2}$, by imposing $\ket{\pm}$ to be eigenstates of $R_{\psi}$, thus giving
|$\ket{\Phi_{\psi}^{\pm}}=2(1-\gamma_{z})\sqrt{1-\lvert\alpha_{-}\rvert^{2}}\ket{+},$ (29)
where |$\ket{\psi}=\alpha_{-}\ket{-}+\sqrt{1-\lvert\alpha_{-}\rvert^{2}}\ket{+}$, with $\alpha_{-}$ that can be chosen to be real because of phase covariance. Such unraveling is shown in Fig. 2 (c). This model shows the flexibility of the generalized RO, which allows us not just to have an effective ensemble consisting of only three states, but also to choose such ensemble in non-unique ways and with different deterministic evolutions for the initial state.
The possibility of controlling the realizations in many different ways is an evident advantage of the new method proposed since it allows us to choose the most convenient ensemble, thus improving the computational efficiency of our model. In Fig. 2 (d), we show the Shannon entropy $H(\{p_{i}\})=-\sum_{i}p_{i}\log_{2}p_{i}$ of the probabilities $\{p_{0},p_{1},p_{\psi}\}$ of the occupations of the states |$\ket{0},\ket{1},\ket{\psi_{\text{det}}(t)}$ for different choices of $\phi_{1}=\lambda\phi_{1}^{\text{lb}}+(1-\lambda)\phi_{1}^{\text{ub}}$. Therefore, it is possible to choose the unravelings in order to minimize the amount of classical information required to describe the average state $\rho=\sum_{i}p_{i}\ket{\psi_{i}}\bra{\psi_{i}}$ |. In Fig. 2 (e), we show that it is possible to have unravelings with vastly different number of jumps involved. This is crucial for the computational efficiency since the fewer jumps involved the more efficient the simulation is, since one does not need to diagonalize the RO at each time-step, but only at the rare times in which a jump happens [63]. This can be done by connecting the jump probability to the reduction of the norm of the deterministic state. This fact is shown in Fig. 2 (f), in which we show a strong dependence between the total number of jumps, with more efficient simulations for fewer jumps.
@@ V.2 Positive unraveling for non-P-divisible dynamics
Figure 3: Positive unravelings for non-P-divisible phase covariant dynamics, $z$ (blue, solid) and $x$ (green, dashed) components of the Bloch vectors. The thick solid lines are the exact results, the dots are obtained with the RO technique. In lighter shade, 5 realizations are shown. (a): $\gamma_{z}<-\frac{1}{2}\sqrt{\gamma_{+}\gamma_{-}}$ as in Eq. (32) with $\kappa=4$; (b): $\gamma_{+}=\gamma_{-}<0$ as in Eq. (35) with $\kappa=0.25$. The unravelings are obtained using $N=10^{4}$ states.
@@ V.2.2 P-divisibility broken by $\gamma_{\pm}$
[…]
and therefore $R$ cannot be positive for these states. On the other hand, by imposing jumps to $\ket{\pm}$, it is possible to have a positive unraveling. Let us focus, for the sake of simplicity, to the case $\gamma_{+}=\gamma_{-}\eqqcolon\gamma$. mposing jumps to $\ket{\pm}$ and solving for |$\ket{\Phi_{\psi}}$ gives us
|$\ket{\Phi_{\psi}^{\pm}}=\frac{\sqrt{1-\lvert\alpha_{-}\rvert^{2}}}{\alpha_{-}}[2\alpha_{-}(\gamma-\gamma_{z})-\phi_{-}]\ket{+}+\phi_{-}\ket{-},$ (34)
@@
Figure 4: Driven dynamics of Eq. (39) for the ratio $\beta/\gamma=10$. (a): $z$ (blue, solid) and $x$ (green, dashed) components of the Bloch vectors. The thick solid lines are the exact results, the dots are obtained with the RO technique. In lighter shade, 5 realizations are shown. The unraveling is obtained using $N=10^{4}$ states. (b): trajectory in the Bloch sphere of $\rho(t)$ (black, dashed) and of |$\ket{\psi_{\text{det}}(t)}$.
@@ V.3 Non-positive unravelings
$\braket{\psi_{\perp}}{R_{\psi,t}}{\psi_{\perp}}=4\lvert\alpha\rvert^{2}(1-\lvert\alpha\rvert^{2})\gamma(t)<0,$ (36)
@@ V.4 Driven dynamics
Except for the constant factor $\gamma$, this is the eternally non-Markovian of Eq. (28), but with the driving term that breaks the phase covariance Eq. (24), since it is not proportional to $\sigma_{z}$.
@@ Appendix A Connection to the master equation
Suppose that, at some time $t$, the state of the system is described by $\rho(t)=\sum_{i}p_{i}\ket{\psi_{i}(t)}\bra{\psi_{i}(t)}$ |. Let us focus on one particular realization |$\ket{\psi_{i}(t)}$. In the infinitesimal time interval $dt$, the state can evolve via a jump
|$\ket{\psi_{i}(t)}\mapsto\ket{\varphi^{j}_{\psi(t),t}},\quad j=1,...,d$ (41)
to an eigenstate of $R_{\psi_{i}(t),t}$, with probability $p^{j}_{\psi_{i}(t),t}=\lambda^{j}_{\psi_{i}(t),t}dt$, with $\lambda^{j}_{\psi_{i}(t),t}$ being the corresponding eigenvalue. Alternatively, it can evolve deterministically
|$\ket{\psi_{i}(t)}\mapsto\ket{\psi^{\text{det}}_{i}(t+dt)}=\frac{(\mathbbm{1}-i\,K_{\psi_{i}(t),t}dt)\ket{\psi_{i}(t)}}{\lVert(\mathbbm{1}-i\,K_{\psi_{i}(t),t}dt)\ket{\psi_{i}(t)}\rVert},$ (42)
@@
$p^{\text{det}}_{\psi_{i}(t),t}=1-\sum_{j}p^{j}_{\psi_{i}(t),t}=1-\operatorname{tr}R_{\psi_{i}(t),t}\,dt.$ (43)
@@
$\lVert(\mathbbm{1}-i\,K_{\psi_{i}(t),t}dt)\ket{\psi_{i}(t)}\rVert^{2}=1-\operatorname{tr}R_{\psi_{i}(t),t}\,dt=p^{\text{det}}_{\psi_{i}(t),t}.$ (44)
@@
|$\begin{split}\ket{\psi_{i}(t)}\bra{\psi_{i}(t)}\mapsto&\,p^{\text{det}}_{\psi_{i}(t),t}\ket{\psi^{\text{det}}_{i}(t+dt)}\bra{\psi^{\text{det}}_{i}(t+dt)}\\ &+dt\,\sum_{j=1}^{d}\lambda^{j}_{\psi_{i}(t),t}\ket{\varphi^{j}_{\psi_{i}(t),t}}\bra{\varphi^{j}_{\psi_{i}(t),t}}\\ =&-i\,dt(K_{\psi_{i}(t),t}\ket{\psi_{i}(t)}\bra{\psi_{i}(t)}-\ket{\psi_{i}(t)}\bra{\psi_{i}(t)}K_{\psi_{i}(t),t}^{\dagger})\\ &+dt\,R_{\psi_{i}(t),t}+\ket{\psi_{i}(t)}\bra{\psi_{i}(t)}.\end{split}$ (45)
Therefore, |$\ket{\psi_{i}(t)}\bra{\psi_{i}(t)}$ | evolves, on average, as by the master equation (1). If one averages over all possible states |$\ket{\psi_{i}(t)}$, the exact master equation (1) is obtained also for $\rho(t)$, thus showing that such unravelings indeed reproduce the exact dynamics.
@@
In the unravelings using the $W$-operator, the jumps are to states orthogonal to the pre-jump state, with positive rates whenever the dynamics is $P$-divisible. Analogous unravelings can also be obtained using $R$ by imposing
$R_{\psi(t),t}\ket{\psi(t)}=\lambda^{\text{det}}\ket{\psi(t)},$ (46)
@@
$C_{\psi(t),t}=c_{11}P_{\psi_{1}}-2\sum_{i=2}^{d}\braket{\psi_{i}}{J_{t}[P_{\psi_{1}}]}{\psi_{1}}\ket{\psi_{i}}\bra{\psi_{1}},$ (47)
where $P_{\psi}=\ket{\psi}\bra{\psi}$ |, with $\operatorname{Re}c_{11}\geq-\braket{J_{t}[P_{\psi(t)}]}_{\psi(t)}$ to ensure the positivity of $\lambda^{\text{det}}$. Besides this constraint, $c_{11}$ is a free parameter that can be used to modify the free evolution. Also, any additional term proportional to |$\ket{\psi_{i}}\bra{\psi_{j}}$ |, $j\geq 2$, in (47) does not affect the unraveling, since $C_{\psi(t),t}$ only acts on |$\ket{\psi(t)}$. Any eigenstate |$\ket{\varphi_{i}}\neq\ket{\psi(t)}$ of $R_{\psi(t),t}$ must be orthogonal to |$\ket{\psi(t)}$, therefore |$\ket{\varphi_{i}}=(\mathbbm{1}-P_{\psi(t)})\ket{\varphi_{i}}$ and
$\begin{split}R_{\psi(t),t}\ket{\varphi_{i}}=&R_{\psi(t),t}(\mathbbm{1}-P_{\psi(t)})\ket{\varphi_{i}}\\ =&\lambda^{i}_{\psi(t),t}\ket{\varphi_{i}}\\ =&\lambda^{i}_{\psi(t),t}(\mathbbm{1}-P_{\psi(t)})\ket{\varphi_{i}}.\end{split}$ (48)
Multiplying on the left by $(\mathbbm{1}-P_{\psi(t)})$, we obtain that |$\ket{\varphi_{i}}$ is an eigenstate also of $W_{\psi}=(\mathbbm{1}-P_{\psi})J_{t}[P_{\psi}](\mathbbm{1}-P_{\psi})$. Therefore, $W_{\psi}$ and $R_{\psi(t),t}^{\perp}$ have the same eigenstates and eigenvalues
$\lambda^{i}_{\psi(t),t}=\sum_{\alpha}\gamma_{\alpha}\lvert\braket{\varphi_{i}}{L_{\alpha}(t)}{\psi(t)}\rvert^{2},$ (49)

### v2

@@ V Control of the realizations
Figure 2: Eternally non-Markovian dynamics, for the initial state |$\ket{\psi(0)}=\alpha\ket{0}+\sqrt{1-\lvert\alpha\rvert^{2}}\ket{1}$, with $\alpha\approx-0.49$. Top: $z$ (blue, solid) and $x$ (green, dashed) components of the Bloch vectors. The thick solid lines are the exact results, the dots are obtained with the RO technique. In lighter shade, 5 realizations are shown. The unravelings are obtained using $N=10^{3}$ states. (a): $\phi_{1}=\phi_{1}^{\text{ub}}$ and only jumps to |$\ket{0}$; (b): $\phi_{1}=\phi_{1}^{\text{lb}}$ and only jumps to |$\ket{1}$; (c): jumps to $\ket{\pm}$. Bottom: characterization of the stochastic realizations for different choices of $\phi_{1}=\lambda\phi_{1}^{\text{lb}}+(1-\lambda)\lambda\phi_{1}^{\text{ub}}$, $0\leq\lambda\leq 1$. (d): Shannon entropy $H(\{p_{i}\})=-\sum_{i}p_{i}\log_{2}p_{i}$ for the probability distribution $\{p_{0},p_{1},p_{\psi}\}$ of the occupations of the states |$\ket{0},\ket{1},\ket{\psi_{\text{det}}(t)}$; (e): number of jumps, using $N=10^{4}$ states; (f): computational time (left axis, solid) and total number of jumps (right axis, dotted).
@@ V.1 Phase covariant dynamics
We now show that, as long as P-divisibility holds, it is always possible to have a positive unraveling using only three states: the eigenstates |$\ket{0},\ket{1}$ of $\sigma_{z}$, and |$\ket{\psi_{\text{det}}(t)}$, the initial state deterministically evolved up to time $t$ according to Eq. (19). The possibility of using such a small effective ensemble drastically simplifies the simulations. The RO is chosen such that |$\ket{\psi_{\text{det}}(t)}$ only jumps to |$\ket{0},\ket{1}$ and, after one jump has occurred, only jumps |$\ket{1}\leftrightarrow\ket{0}$ are present, without any deterministic evolution. Thus, the effective ensemble used for the simulations only contains three states. The possibility of considering such finite (and small) effective ensemble, significantly enhances the computational efficiency of this method, since one is not required to compute at each time-step the evolution of all states, but only needs to update the probability of occupation of such states. This fact is particularly interesting in comparison with the NMQJ method that, expect for some special cases, needs infinitely many states in the effective ensemble.
If no jumps have occurred, it is possible to have jumps |$\ket{\psi(t)}\mapsto\ket{0},\ket{1}$, with |$\ket{\psi(t)}$ deterministically evolved according to Eq. (19), by imposing that |$\ket{0}$ is an eigenstate of $\Psi\text{-}R_{\psi(t)}$. This corresponds to a transformation defining the RO of Eq. (17) of the form
|$\ket{\Phi_{{\psi(t)}}}=\alpha\left(2\gamma_{z}-\frac{\phi_{1}}{\sqrt{1-\lvert\alpha\rvert^{2}}}\right)\ket{0}+\phi_{1}\ket{1},$ (27)
where |$\ket{\psi(t)}=\alpha\ket{0}+\sqrt{1-\lvert\alpha\rvert^{2}}\ket{1}$, with $\alpha$ that, without loss of generality, can be assumed to be real because of phase covariance, while $\phi_{1}$ which can be chosen freely inside a suitable time- and state-dependent interval $\phi_{1}\in[\phi_{1}^{\text{lb}},\phi_{1}^{\text{ub}}]$ (for the details, the definition of $\phi_{1}^{\text{lb}},\phi_{1}^{\text{ub}}$ and the proof of the positivity, see Appendix C). The freedom in choosing $\phi_{1}$ allows us to have different realizations for the unraveling: even if the post-jump states are fixed, it is possible to unravel the dynamics with different jump rates and deterministic evolutions. In particular, for $\phi_{1}=\phi_{1}^{\text{lb}}$, only jumps |$\ket{\psi(t)}\mapsto\ket{1}$ are allowed, while for $\phi_{1}=\phi_{1}^{\text{ub}}$, only |$\ket{\psi(t)}\mapsto\ket{0}$. On the other hand, for any $\phi_{1}=\lambda\phi_{1}^{\text{lb}}+(1-\lambda)\phi_{1}^{\text{ub}}$, $10<\lambda<$ 1, the unraveling remains positive, with jumps to both eigenstates of $\sigma_{z}$ and with the possibility of choosing different jump rates depending on $\lambda$.
@@ V.1.1 Eternally non-Markovian dynamics
The negativity of $\gamma_{z}$ at all times implies that such dynamics is CP-indivisible at all times, and therefore cannot be unraveled using the standard MCWF nor its non-Markovian generalization [30, 29] because of the negativity of the rate since the very beginning of time. However, it is possible to unravel it using the generalized RO. In addition, it is possible to realize qualitatively different realizations of the stochastic process. In Fig. 2 (a)-(b), we show unravelings obtained either with only jumps to |$\ket{0}$ or to |$\ket{1}$ by suitably choosing $\phi_{1}$. The code used for obtaining the simulations is available at [62]. Additionally, it can also be unraveled with jumps to |±=(|$\ket{\pm}=(\ket{0}\pm\ket{1})/\sqrt{2}$, by imposing $\ket{\pm}$ to be eigenstates of $\Psi\text{-}R_{\psi(t)}$, thus giving
|$\ket{\Phi_{{\psi(t)}}^{\pm}}=2(1-\gamma_{z})\sqrt{1-\lvert\alpha_{-}\rvert^{2}}\ket{+},$ (29)
where |$\ket{\psi(t)}=\alpha_{-}\ket{-}+\sqrt{1-\lvert\alpha_{-}\rvert^{2}}\ket{+}$, with $\alpha_{-}$ that can be chosen to be real because of phase covariance. Such unraveling is shown in Fig. 2 (c). This model shows the flexibility of the generalized RO, which allows us not just to have an effective ensemble consisting of only three states, but also to choose such ensemble in non-unique ways and with different deterministic evolutions for the initial state.
The possibility of controlling the realizations in many different ways is an evident advantage of the new method proposed since it allows us to choose the most convenient ensemble, thus improving the computational efficiency of our model. In Fig. 2 (d), we show the Shannon entropy $H(\{p_{i}\})=-\sum_{i}p_{i}\log_{2}p_{i}$ of the probabilities $\{p_{0},p_{1},p_{\psi}\}$ of the occupations of the states |$\ket{0},\ket{1},\ket{\psi_{\text{det}}(t)}$ for different choices of $\phi_{1}=\lambda\phi_{1}^{\text{lb}}+(1-\lambda)\phi_{1}^{\text{ub}}$. Therefore, it is possible to choose the unravelings in order to minimize the amount of classical information required to describe the average state $\rho=\sum_{i}p_{i}{P_{\psi_{i}}}$. In Fig. 2 (e), we show that it is possible to have unravelings with vastly different number of jumps involved. This is crucial for the computational efficiency since the fewer jumps involved the more efficient the simulation is, since one does not need to diagonalize the RO at each time-step, but only at the rare times in which a jump happens [63]. This can be done by connecting the jump probability to the reduction of the norm of the deterministic state. This fact is shown in Fig. 2 (f), in which we show a strong dependence between the total number of jumps, with more efficient simulations for fewer jumps.
@@ V.2 Positive unraveling for non-P-divisible dynamics
Figure 3: Positive unravelings for non-P-divisible phase covariant dynamics for the initial state |$\ket{\psi(0)}=\alpha\ket{0}+\sqrt{1-\lvert\alpha\rvert^{2}}\ket{1}$, with $\alpha\approx 0.92$. $z$ (blue, solid) and $x$ (green, dashed) components of the Bloch vectors. The thick solid lines are the exact results, the dots are obtained with the RO technique. In lighter shade, 5 realizations are shown. (a): $\gamma_{z}<-\frac{1}{2}\sqrt{\gamma_{+}\gamma_{-}}$ as in Eq. (32) with $\kappa=4$; (b): $\gamma_{+}=\gamma_{-}<0$ as in Eq. (35) with $\kappa=0.25$. The unravelings are obtained using $N=10^{4}$ states.
@@ V.2.2 P-divisibility broken by $\gamma_{\pm}$
$\braket{0}{\mathcal{J}_{t}[{P_{1}}]}{0}=\gamma_{-}<0,$ (33)
and therefore $\Psi\text{-}R$ cannot be positive for these states. On the other hand, by imposing jumps to $\ket{\pm}$, it is possible to have a positive unraveling. Let us focus, for the sake of simplicity, to the case $\gamma_{+}=\gamma_{-}\eqqcolon\gamma$. Imposing jumps to $\ket{\pm}$ and solving for |$\ket{\Phi_{{\psi(t)}}}$ gives us
|$\ket{\Phi_{{\psi(t)}}^{\pm}}=\frac{\sqrt{1-\lvert\alpha_{-}\rvert^{2}}}{\alpha_{-}}[2\alpha_{-}(\gamma-\gamma_{z})-\phi_{-}]\ket{+}+\phi_{-}\ket{-},$ (34)
@@
Figure 4: Driven dynamics of Eq. (39) for the ratio $\beta/\gamma=10$ and initial state |$\ket{\psi(0)}=\alpha\ket{0}+\sqrt{1-\lvert\alpha\rvert^{2}}\ket{1}$, with $\alpha\approx 0.20$. (a): $z$ (blue, solid) and $x$ (green, dashed) components of the Bloch vectors. The thick solid lines are the exact results, the dots are obtained with the RO technique. In lighter shade, 5 realizations are shown. The unraveling is obtained using $N=10^{4}$ states. (b): trajectory in the Bloch sphere of $\rho(t)$ (black, dashed) and of |$\ket{\psi_{\text{det}}(t)}$.
@@ V.3 Non-positive unravelings
$\braket{\psi_{\perp}}{{\Psi\text{-}R_{\psi}}}{\psi_{\perp}}=4\lvert\alpha\rvert^{2}(1-\lvert\alpha\rvert^{2})\gamma(t)<0,$ (36)
@@ V.4 Driven dynamics
The jump term, except for the constant factor $\gamma$, is the same as the eternally non-Markovian of Eq. (28). However, the dynamics is not phase covariant since the driving, not being proportional to $\sigma_{z}$, breaks phase covariance Eq. (24).
@@ Appendix A Connection to the master equation
Suppose that, at some time $t$, the state of the system is described by $\rho(t)=\sum_{i}p_{i}{P_{\psi_{i}(t)}}$. Let us focus on one particular realization |$\ket{\psi_{i}(t)}$. In the infinitesimal time interval $dt$, the state can evolve via a jump
|$\ket{\psi_{i}(t)}\mapsto{\ket{\varphi^{j}_{\psi(t)}}},\quad j=1,...,d$ (41)
to an eigenstate of $\Psi\text{-}R_{\psi_{i}(t)}$, with probability $p^{j}_{\psi_{i}(t)}=\lambda^{j}_{\psi_{i}(t)}dt$, with $\lambda^{j}_{\psi_{i}(t)}$ being the corresponding eigenvalue. Alternatively, it can evolve deterministically
|$\ket{\psi_{i}(t)}\mapsto\ket{\psi^{\text{det}}_{i}(t+dt)}={\frac{(\mathbbm{1}-i\,K_{\psi_{i}(t)}dt)\ket{\psi_{i}(t)}}{\lVert(\mathbbm{1}-i\,K_{\psi_{i}(t)}dt)\ket{\psi_{i}(t)}\rVert}},$ (42)
@@
${p^{\text{det}}_{\psi_{i}(t)}=1-\sum_{j}p^{j}_{\psi_{i}(t)}=1-\operatorname{tr}\left[\Psi\text{-}R_{\psi_{i}(t)}\right]\,dt.}$ (43)
@@
${\lVert(\mathbbm{1}-i\,K_{\psi_{i}(t)}dt)\ket{\psi_{i}(t)}\rVert^{2}=1-\operatorname{tr}\left[\Psi\text{-}R_{\psi_{i}(t)}\right]\,dt=p^{\text{det}}_{\psi_{i}(t)}}.$ (44)
@@
$\begin{split}{P_{\psi_{i}(t)}}\mapsto&\,{p^{\text{det}}_{\psi_{i}(t)}}{P_{\psi^{\text{det}}_{i}(t+dt)}}\\ &+dt\,\sum_{j=1}^{d}{\lambda^{j}_{\psi_{i}(t)}{P_{\varphi^{j}_{\psi_{i}(t)}}}}\\ =&-i\,dt{(K_{\psi_{i}(t)}{P_{\psi_{i}(t)}}-{P_{\psi_{i}(t)}}K_{\psi_{i}(t)}^{\dagger})}\\ &+dt\,{\Psi\text{-}R_{\psi_{i}(t)}+{P_{\psi_{i}(t)}}}.\end{split}$ (45)
Therefore, ${P_{\psi_{i}(t)}}$ evolves, on average, as by the master equation (1). If one averages over all possible states |$\ket{\psi_{i}(t)}$, the exact master equation (1) is obtained also for $\rho(t)$, thus showing that such unravelings indeed reproduce the exact dynamics.
@@
In the unravelings using the $W$-operator, the jumps are to states orthogonal to the pre-jump state, with positive rates whenever the dynamics is $P$-divisible. Analogous unravelings can also be obtained using $\Psi\text{-}R$ by imposing
${\Psi\text{-}R_{\psi(t)}}\ket{\psi(t)}=\lambda^{\text{det}}\ket{\psi(t)},$ (46)
@@
${C_{\psi(t)}}=c_{11}P_{\psi_{1}}-2\sum_{i=2}^{d}\braket{\psi_{i}}{{\mathcal{J}}_{t}[P_{\psi_{1}}]}{\psi_{1}}\ket{\psi_{i}}\bra{\psi_{1}},$ (47)
with $\operatorname{Re}c_{11}\geq-\braket{{\mathcal{J}}_{t}[P_{\psi(t)}]}_{\psi(t)}$ to ensure the positivity of $\lambda^{\text{det}}$. Besides this constraint, $c_{11}$ is a free parameter that can be used to modify the free evolution. Also, any additional term proportional to |$\ket{\psi_{i}}\bra{\psi_{j}}$ |, $j\geq 2$, in (47) does not affect the unraveling, since ${C_{\psi(t)}}$ only acts on |$\ket{\psi(t)}$. Any eigenstate |$\ket{\varphi_{i}}\neq\ket{\psi(t)}$ of $\Psi\text{-}R_{\psi(t)}$ must be orthogonal to |$\ket{\psi(t)}$, therefore |$\ket{\varphi_{i}}=(\mathbbm{1}-P_{\psi(t)})\ket{\varphi_{i}}$ and
$\begin{split}{\Psi\text{-}R_{\psi(t)}}\ket{\varphi_{i}}=&{\Psi\text{-}R_{\psi(t)}}(\mathbbm{1}-P_{\psi(t)})\ket{\varphi_{i}}\\ =&{\lambda^{i}_{\psi(t)}}\ket{\varphi_{i}}\\ =&{\lambda^{i}_{\psi(t)}}(\mathbbm{1}-P_{\psi(t)})\ket{\varphi_{i}}.\end{split}$ (48)
Multiplying on the left by $(\mathbbm{1}-P_{\psi(t)})$, we obtain that |$\ket{\varphi_{i}}$ is an eigenstate also of $W_{\psi}=(\mathbbm{1}-P_{\psi}){\mathcal{J}}_{t}[P_{\psi}](\mathbbm{1}-P_{\psi})$. Therefore, $W_{\psi}$ and $\Psi\text{-}R_{\psi(t)}$ have the same eigenstates and eigenvalues
${\lambda^{i}_{\psi(t)}}=\sum_{\alpha}\gamma_{\alpha}\lvert\braket{\varphi_{i}}{L_{\alpha}(t)}{\psi(t)}\rvert^{2},$ (49)